# Functional Graphical Models: Structure Enables Offline Data-Driven Optimization

### v2

@@ 1 Introduction
Machine learning models are typically trained for prediction, but we often want to use such models to solve optimization or decision-making problems. Imagine we would like to use a dataset of fluorescent proteins and their fluorescence levels to design a new protein with the highest possible fluorescence (Brookes et al., 2019; Trabucco et al., 2022). Similarly, when provided with examples of hardware accelerators and their performance, an engineer might want to infer a more performant hardware accelerator design (Kao et al., 2020; Kumar et al., 2021). A direct machine learning approach to such problems would be to train a surrogate model that predicts the performance for a given design, and then find its maximizers with some optimization method (Trabucco et al., 2021). In this work, we focus on the challenging yet practical scenario where we can use offline training data to train a surrogate model, but cannot collect additional samples online. This offline data-driven optimization (DDO) approach is often applied in model-based optimization (Hoburg & Abbeel, 2014; Gómez-Bombarelli et al., 2018; Trabucco et al., 2021, MBO) and offline reinforcement learning (Levine et al., 2020; Fujimoto & Gu, 2021, RL).
@@ 2 Related Work
The idea of using data and machine learning tools for design and optimization has been one of the primary motivations of Bayesian optimization (BO) (Brochu et al., 2010; Shahriari et al., 2015). In BO, one establishes a prior belief of the considered function and updates it based on the given dataset, as well as queries for function evaluations that balance maximization and exploration of the function. However, such an online approach is not applicable to problems where additional queries cannot be made or are very cost- and time-consuming (e.g., drug design) that would ideally be solved fully offline. Nevertheless, we consider our work to be of interest to the BO community, where it is common to assume a decomposition of the objective over a dependency graph, which simplifies the problem and enables solving it with fewer queries (Kandasamy et al., 2015; Rolland et al., 2018; Hoang et al., 2018; Ziomek & Bou-Ammar, 2023). In Section 4, we prove that such a decomposition exists for every function satisfying mild conditions, enabling tractable optimization with finite amounts of offline data.
@@
There has been a growing interest in model-based optimization (MBO) and the specific instantiation of MBO from static data, or data-driven optimization (DDO). Recent works have investigated computational design via such offline MBO for chemical molecules, proteins, hardware accelerators, and more (Gómez-Bombarelli et al., 2018; Brookes et al., 2019; Kumar & Levine, 2020; Kumar et al., 2021). MBO and DDO are often used interchangeably, and we employ the term DDO to refer specifically to optimization from static offline data. Most of these prior methods adopt a “conservative” (i.e., pessimistic) approach to DDO, employing regularizers to prevent the generation of out-of-distribution designs (Trabucco et al., 2021; Qi et al., 2022). Combined with appropriate inductive bias, these methods found application in practical engineering problems, such as hardware accelerator (microchip) design (Kumar et al., 2021). In our paper, we instead investigate how discovering functional independence structure via our proposed FGM framework can enable efficient optimization from data, and provide to our knowledge the first theoretical analysis of sample efficiency for DDO with such independence properties.
@@ 3 Preliminaries
where $\Pi$ consists of probability distributions over $\mathcal{X}$, and $\mathcal{F}$ is a regression class that consists of functions mapping $\mathcal{X}$ to $\mathbb{R}$. When $\Pi$ is a fully expressive class [^1] When $\mathcal{X}$ is discrete, we can always use this fully expressive class. However, when $\mathcal{X}$ is continuous, as we will see in Appendix  B.3, the coverage term could be infinite when $\Pi$ consists of delta distributions. For example, we might use $\Pi=\{\mathcal{N}({\mathbf{x}},\sigma^{2});{\mathbf{x}}\in\mathcal{X}$} for small $\sigma^{2}$ in the continuous setting., i.e., $\Pi=\Delta(\mathcal{X})$, Step (b) reduces to
@@
The primary practical challenge in the fully offline DDO setting lies in dealing with insufficient data coverage. To see this, consider the naïve approach: we can set a model $\hat{f}({\mathbf{x}})$ to match the observed $f({\mathbf{x}})$ values and assign random values to examples outside of the offline data. This approach amounts to constructing a surrogate model $\hat{f}$ in the initial Step (a) without relying on any structural bias. In the discrete case, it is equivalent to setting =$\mathcal{F}=\{{\mathbf{x}}\mapsto{\bm{\theta}}^{\top}{\bm{\phi}}({\mathbf{x}})\ |\ {\bm{\theta}}\in\mathbb{R}^{|\mathcal{X}|}\}$, where $\phi({\mathbf{x}})$ is a $|\mathcal{X}|$-dimensional one-hot encoding vector. Maximizing such a model corresponds to choosing ${\mathbf{x}}$ for which $\theta_{{\mathbf{x}}}$ is highest. However, if the input space $\mathcal{X}$ is not fully covered by the data, such an optimization process may end up outputting a novel design ${\mathbf{x}}$ whose ground-truth value $f({\mathbf{x}})$ is very low, but its random prediction $\hat{f}({\mathbf{x}})$ happens to be high, thus adversarially exploiting the model’s errors. One may want to tackle this problem by, instead of assigning random values to unobserved designs, modeling them pessimistically—for example, by setting them to $\hat{f}({\mathbf{x}})=-\infty$. This leads to selecting the following input:
@@ 4.1 Functional Independence
If $S=\emptyset$ and there exist mutually disjoint supersets $\underline{A}$ of $A$ and $\underline{B}$ of $B$ such that $f_{-B}({\mathbf{x}}_{-B})=f_{\underline{A}}({\mathbf{x}}_{\underline{A}})$, and $f_{-A}({\mathbf{x}}_{-A})=f_{\underline{B}}({\mathbf{x}}_{\underline{B}})$ then the independence is absolute.
@@ 4.2 FGMs
Our construction of FGMs is analogous to that of probabilistic graphical models (PGMs), where the absence of an edge indicates probabilistic conditional independence between two random variables (Jordan, 1999; Koller & Friedman, 2009). As we prove in Appendix B.1, a result similar to the Hammersley-Clifford theorem in PGMs (Clifford & Hammersley, 1971) holds for FGMs, decomposing arbitrary functions as follows.
@@ 5.1 Regret Guarantees
First, we review the regret guarantee of the naïve approach in Step (a) and Step (b). A standard regret guarantee in this setting (Chen & Jiang, 2019) states that, with probability $1-\delta$, the following holds:
@@
where $p({\mathbf{x}}_{C}),\bar{\pi}({\mathbf{x}}_{c})$ are, with slight abuse of notation, marginal distributions of ${\mathbf{x}}_{C}$ with respect to $p({\mathbf{x}})$ and $\pi({\mathbf{x}})$.
While the bound of the naïve approach in Eq. (3) implies that MBO without FGM structure necessitates the data distribution $p({\mathbf{x}})$ to cover the optimal design itself, the FGM factorization only necessitates the data distribution to cover components within each clique in the optimal design. [^3] Technically, Theorem 2 states that all distributions in the class $\Pi$ need to be covered, not just the best distribution $\pi^{\star}$, but . As we will mention in Section 5.3, $\max_{\bar{\pi}\in\Pi}$ can be replaced with $\pi^{\star}$ by incorporating conservative approaches. This is a significantly less stringent requirement, as the best design in the offline data might be much less effective than the best design achievable by combining favorable settings of the values in each of the cliques, as illustrated in Figure 2. We will delve further into the implications of this below.
@@
Note that the literature on additive models (Hastie & Tibshirani, 1987) has previously explored their role in reducing statistical complexity. However, our innovation lies in characterizing the coverage term in the context of DDO. It’s important to highlight that simply reducing statistical complexity polynomially (i.e., improving the term $(a^{\prime})$) is insufficient for overcoming the curse of dimensionality, because the term $(a)$ might still grow exponentially.
@@ 6 Graph Discovery
This assumption may appear strong. However, whitening to remove correlations is a common procedure in machine learning, and in cases where the data distribution has higher order moments, we can use more advanced representation learning methods, such as variational auto-encoders (Kingma & Welling, 2013; Rezende & Mohamed, 2015; Dai & Wipf, 2019, VAE). Indeed, these methods are designed specifically to acquire latent representations of the data that follow the standard Gaussian distribution.
@@ 7 Experiments
In this section, we present numerical experiments to verify our theoretical findings as well as to examine how readily applicable they are in practice. We start with the setting where the FGM is known, and then explore FGM discovery methods, including with learned latent spaces.
@@
Since we do not assume access to the function’s FGM, nor Gaussianity of the data, we begin by learning marginally-Gaussian representations ${\mathbf{z}}({\mathbf{x}}_{\text{obs}})$ of the designs. In our experiments, we choose to learn them with Variational Auto-Encoders (Kingma & Welling, 2013, VAE). Viewing the design values as a function of the Gaussian representations, ${\textnormal{y}}=f({\mathbf{z}})$, we then compute the estimator from Eq. (6) to infer the FGM of $f({\mathbf{z}})$ and extract its clique set $\hat{\mathcal{C}}$. We then learn a predictive neural network model that decomposes over these cliques, and perform gradient ascent with respect to it over the ${\mathbf{z}}$ variable. Then, we decode the representation of such a design $\hat{{\mathbf{z}}}$ with our VAE decoder to propose a new design candidate $\hat{{\mathbf{x}}}_{\text{obs}}$. We summarize the entire algorithm for DDO with VAEs and FGMs in Algorithm 1.
@@
We compare this method with naïve gradient ascent (GA), reward-weighted regression (Peters & Schaal, 2007, RWR)—to see the effectiveness of naïve methods in the tasks we designed; conservative objective models (Trabucco et al., 2021, COMs)—to compare to an off-the-shelf state-of-the-art method; and an ablative method that learns VAE representations and performs gradient ascent directly on them (VAE-GA), similarly to Gómez-Bombarelli et al. (2018)—to isolate the benefits of FGM decomposition on top of the learned latent space.
@@
One may wonder if Gaussianity of both the distribution of the ground-truth inputs ${\mathbf{x}}$ and the VAE prior simplifies the task for our algorithm, since the VAE preprocessing step could learn to invert our non-linear data transformation. This, however, is very unlikely given that there are infinitely many bijective reparameterizations of ${\mathbf{x}}$ that could explain $f({\mathbf{x}})$ perfectly and be also Gaussian (e.g., rotations of ${\mathbf{x}}$). Nevertheless, as a sanity check, we re-ran the experiments with the ground-truth inputs ${\mathbf{x}}$ sampled from a mixture of Gaussians, rather than the unimodal Gaussian distribution. This way we made sure that the representations learned by the VAE are significantly distinct from the ground-truth inputs, as well as made the multi-modality of the observed designs ${\mathbf{x}}_{\text{obs}}$ more explicit. Figure 5 shows that our findings are preserved under this setting, suggesting that the role of the VAE is confined to reparameterization of the designs.
@@ Appendix A Auxiliary Facts
Theorem 3 (Hammersley-Clifford (Clifford & Hammersley, 1971)).
@@ B.2 Proof of Theorem 1
That is, measure $e_{f}({\mathbf{x}})$ factorizes over $\{\mathcal{V}\setminus\{i\},\mathcal{V}\setminus\{j\}\}$. Hence, random variables ${\textnormal{x}}_{i}$ and ${\textnormal{x}}_{j}$ are (probabilistically) independent given ${\mathbf{x}}_{\mathcal{V}\setminus\{i,j\}}$ (Studeny, 2006). As the pair $(i,j)$ was arbitrary, this implies that $e_{f}({\mathbf{x}})$ satisfies the pairwise Markov property with respect to $\mathcal{G}$ (equivalently, $\mathcal{G}$ is an I-map of $e_{f}({\mathbf{x}})$). As the measure induced by $e_{f}({\mathbf{x}})$ is strictly positive, Hammersley-Clifford Theorem (Clifford & Hammersley, 1971) implies that $e_{f}({\mathbf{x}})$ factorizes according to $\mathcal{G}$. That is, for the clique set $\mathcal{C}$ of $\mathcal{G}$, there exist positive functions $\{u_{C}({\mathbf{x}}_{C})\ |\ C\in\mathcal{C}\}$, such that
@@ B.3 Proof of Theorem 2
Following the standard literature of regression (Agarwal et al., 2019) and using Assumption (3), by leveraging $f_{C}\in\mathcal{F}_{C}$, we can obtain with probability $1-\delta$,
@@
$\geq\sum_{C}\mathbb{E}_{{\mathbf{x}}\sim p}\left[\{\hat{f}_{C}({\mathbf{x}}_{C})-f_{C}({\mathbf{x}}_{C})\}^{2}\right]+\sigma\times\left\{\mathbb{E}_{{\mathbf{x}}\sim p}\left[\{\hat{f}_{C_{1}}({\mathbf{x}}_{C_{1}})-f_{C_{1}}({\mathbf{x}}_{C_{1}})\}^{2}\right]^{1/2}\right\}\times\left\{\mathbb{E}_{{\mathbf{x}}\sim p}\left[\left\{\hat{f}_{C_{2}}({\mathbf{x}}_{C_{2}})-f_{C_{2}}({\mathbf{x}}_{C_{2}})\right\}^{2}\right]^{1/2}\right\}$ (11)

### v3

@@ 1 INTRODUCTION
Machine learning models are typically trained for prediction, but we often want to use such models to solve optimization or decision-making problems. Imagine we would like to use a dataset of fluorescent proteins and their fluorescence levels to design a new protein with the highest possible fluorescence (Brookes et al., 2019; Trabucco et al., 2022). Similarly, when provided with examples of hardware accelerators and their performance, an engineer might want to infer a more performant hardware accelerator design (Kao et al., 2020; Kumar et al., 2021). A direct machine learning approach to such problems would be to train a surrogate model that predicts the performance for a given design, and then find its maximizers with some optimization method (Trabucco et al., 2021). In this work, we focus on the challenging yet practical scenario where we can use offline training data to train a surrogate model, but cannot collect additional samples online. This offline data-driven optimization (DDO) approach is often applied in model-based optimization (Hoburg and Abbeel, 2014; Gómez-Bombarelli et al., 2018; Trabucco et al., 2021, MBO) and offline reinforcement learning (Levine et al., 2020; Fujimoto and Gu, 2021, RL).
@@ 2 RELATED WORK
The idea of using data and machine learning tools for design and optimization has been one of the primary motivations of Bayesian optimization (BO) (Brochu et al., 2010; Shahriari et al., 2015). In BO, one establishes a prior belief of the considered function and updates it based on the given dataset, as well as queries for function evaluations that balance maximization and exploration of the function. However, such an online approach is not applicable to problems where additional queries cannot be made or are very cost- and time-consuming (e.g., drug design) that would ideally be solved fully offline. Nevertheless, we consider our work to be of interest to the BO community, where it is common to assume a decomposition of the objective over a dependency graph, which simplifies the problem and enables solving it with fewer queries (Kandasamy et al., 2015; Rolland et al., 2018; Hoang et al., 2018; Ziomek and Bou-Ammar, 2023). In Section 4, we existence of such a decomposition for every function under mild conditions, enabling tractable optimization with finite amounts of offline data.
@@
There has been a growing interest in model-based optimization (MBO) and the specific instantiation of MBO from static data, or data-driven optimization (DDO). Recent works have investigated computational design via such offline MBO for chemical molecules, proteins, hardware accelerators, and more (Gómez-Bombarelli et al., 2018; Brookes et al., 2019; Kumar and Levine, 2020; Kumar et al., 2021). MBO and DDO are often used interchangeably, and we employ the term DDO to refer specifically to optimization from static offline data. Most of these prior methods adopt a “conservative” (i.e., pessimistic) approach to DDO, employing regularizers to prevent the generation of out-of-distribution designs (Trabucco et al., 2021; Qi et al., 2022). Combined with appropriate inductive bias, these methods found application in practical engineering problems, such as hardware accelerator (microchip) design (Kumar et al., 2021). In our paper, we instead investigate how discovering functional independence structure via our FGM framework can enable efficient optimization from data, and provide to our knowledge the first theoretical analysis of sample efficiency for DDO with such independence properties.
@@ 3 PRELIMINARIES
where $\Pi$ consists of probability distributions over $\mathcal{X}$, and $\mathcal{F}$ is a regression class that consists of functions mapping $\mathcal{X}$ to $\mathbb{R}$. When $\Pi$ is a fully expressive class [^1] When $\mathcal{X}$ is discrete, we can always use this fully expressive class. However, when $\mathcal{X}$ is continuous, one must use a restricted distribution class, e.g., $\Pi=\{\mathcal{N}({\mathbf{x}},\sigma^{2});{\mathbf{x}}\in\mathcal{X}$} for small $\sigma^{2}$., i.e., $\Pi=\Delta(\mathcal{X})$, Step (b) reduces to
@@
The primary practical challenge in the fully offline DDO setting lies in dealing with insufficient data coverage. To see this, consider the naïve approach: we can set a model $\hat{f}({\mathbf{x}})$ to match the observed $f({\mathbf{x}})$ values and assign random values to examples outside of the offline data. This approach amounts to constructing a surrogate model $\hat{f}$ in the initial Step (a) without relying on any structural bias. In the discrete case, it is equivalent to setting =$\mathcal{F}=\{{\mathbf{x}}\mapsto{\bm{\theta}}^{\top}{\bm{\phi}}({\mathbf{x}})\ |\ {\bm{\theta}}\in\mathbb{R}^{|\mathcal{X}|}\}$, where $\phi({\mathbf{x}})$ is a $|\mathcal{X}|$-dimensional one-hot encoding vector. Maximizing such a model corresponds to choosing ${\mathbf{x}}$ for which $\theta_{{\mathbf{x}}}$ is highest. However, if the input space $\mathcal{X}$ is not fully covered by the data, such an optimization process may end up outputting a novel design ${\mathbf{x}}$ whose ground-truth value $f({\mathbf{x}})$ is very low, but its random prediction $\hat{f}({\mathbf{x}})$ happens to be high, thus adversarially exploiting the model’s errors. One may want to tackle this problem by, instead of assigning random values to unobserved designs, modeling them pessimistically—for example, by setting them to $\hat{f}({\mathbf{x}})=-\infty$. The resulting solution is:
@@ 4.1 Functional Independence
If $S=\emptyset$ and there exist mutually disjoint supersets $\underline{A}$ of $A$ and $\underline{B}$ of $B$ such that $f_{-B}({\mathbf{x}}_{-B})=f_{\underline{A}}({\mathbf{x}}_{\underline{A}})$, and $f_{-A}({\mathbf{x}}_{-A})=f_{\underline{B}}({\mathbf{x}}_{\underline{B}})$ the independence is absolute.
@@ 4.2 FGMs
Our construction of FGMs is analogous to that of probabilistic graphical models (PGMs), where the absence of an edge indicates probabilistic conditional independence between two random variables (Jordan, 1999; Koller and Friedman, 2009). As we prove in Appendix B.1, a result similar to the Hammersley-Clifford theorem in PGMs (Clifford and Hammersley, 1971) holds for FGMs, decomposing arbitrary functions as follows.
@@ 5.1 REGRET GUARANTEES
First, we review the regret guarantee of the naïve approach in Step (a) and Step (b). A standard regret guarantee in this setting (Chen and Jiang, 2019) states that, with probability $1-\delta$, the following holds:
@@
where $p({\mathbf{x}}_{C}),\bar{\pi}({\mathbf{x}}_{c})$ denote the marginal distributions of ${\mathbf{x}}_{C}$ with respect to $p({\mathbf{x}})$ and $\pi({\mathbf{x}})$.
While the bound of the naïve approach in Eq. (3) implies that MBO without FGM structure necessitates the data distribution $p({\mathbf{x}})$ to cover the optimal design itself, the FGM factorization only necessitates the data distribution to cover components within each clique in the optimal design. This is a significantly less stringent requirement, as the best design in the offline data might be much less effective than the best design achievable by combining favorable settings of the values in each of the cliques, as illustrated in Figure 2, and analyzed deeper below.
@@
Note that the literature on additive models (Hastie and Tibshirani, 1987) has previously explored their role in reducing statistical complexity. In contrast, we characterize the coverage term in the context of DDO. It’s important to highlight that simply reducing statistical complexity polynomially (i.e., improving the term $(a^{\prime})$) is insufficient for overcoming the curse of dimensionality, because the term $(a)$ might still grow exponentially.
@@ 6 GRAPH DISCOVERY
This assumption may appear strong. However, whitening to remove correlations is a common procedure in machine learning, and in cases where the data distribution has higher order moments, we can use more advanced representation learning methods, such as variational auto-encoders (Kingma and Welling, 2013; Rezende and Mohamed, 2015; Dai and Wipf, 2019, VAE). Indeed, these methods are designed specifically to acquire latent representations of the data that follow the standard Gaussian distribution.
@@ 7 EXPERIMENTS
In this section, we present numerical experiments to verify our theoretical findings and examine how readily applicable they are in practice. We start with the setting where the FGM is known, and then explore FGM discovery methods with learned latent spaces.
@@
Since we do not assume access to the function’s FGM, nor Gaussianity of the data, we begin by learning marginally-Gaussian representations ${\mathbf{z}}({\mathbf{x}}_{\text{obs}})$ of the designs. In our experiments, we choose to learn them with Variational Auto-Encoders (Kingma and Welling, 2013, VAE). Viewing the design values as a function of the Gaussian representations, ${\textnormal{y}}=f({\mathbf{z}})$, we then compute the estimator from Eq. (6) to infer the FGM of $f({\mathbf{z}})$ and extract its clique set $\hat{\mathcal{C}}$. We then learn a predictive neural network model that decomposes over these cliques, and perform gradient ascent with respect to it over the ${\mathbf{z}}$ variable. Then, we decode the representation of such a design $\hat{{\mathbf{z}}}$ with our VAE decoder to propose a new design candidate $\hat{{\mathbf{x}}}_{\text{obs}}$. We summarize the entire algorithm for DDO with FGMs in Algorithm 1.
@@
We compare this method with naïve gradient ascent (GA), reward-weighted regression (Peters and Schaal, 2007, RWR)—to see the effectiveness of naïve methods in the tasks we designed; conservative objective models (Trabucco et al., 2021, COMs)—to compare to an off-the-shelf state-of-the-art method; and an ablative method that learns VAE representations and performs gradient ascent directly on them (VAE-GA), similarly to (Gómez-Bombarelli et al., 2018)—to isolate the benefits of FGM decomposition on top of the learned latent space.
@@
One may wonder if Gaussianity of both the distribution of the ground-truth inputs ${\mathbf{x}}$ and the VAE prior simplifies the task for our algorithm, since the VAE preprocessing step could learn to invert our non-linear data transformation. This, however, is very unlikely given that there are infinitely many bijective reparameterizations of ${\mathbf{x}}$ that could explain $f({\mathbf{x}})$ perfectly and be also Gaussian (e.g., rotations of ${\mathbf{x}}$). Nevertheless, as a sanity check, we re-ran the experiments with the ground-truth inputs ${\mathbf{x}}$ sampled from a mixture of Gaussians, rather than the unimodal Gaussian distribution. This way we made sure that the representations learned by the VAE are significantly distinct from the ground-truth inputs, as well as made the multi-modality of the observed designs ${\mathbf{x}}_{\text{obs}}$ more explicit. Figure 5 shows that our findings are preserved under this setting.
@@ Appendix A Auxiliary Facts
Theorem 3 (Hammersley-Clifford (Clifford and Hammersley, 1971)).
@@ B.2 Proof of Theorem 1
That is, measure $e_{f}({\mathbf{x}})$ factorizes over $\{\mathcal{V}\setminus\{i\},\mathcal{V}\setminus\{j\}\}$. Hence, random variables ${\textnormal{x}}_{i}$ and ${\textnormal{x}}_{j}$ are (probabilistically) independent given ${\mathbf{x}}_{\mathcal{V}\setminus\{i,j\}}$ studeny2006probabilistic. As the pair $(i,j)$ was arbitrary, this implies that $e_{f}({\mathbf{x}})$ satisfies the pairwise Markov property with respect to $\mathcal{G}$ (equivalently, $\mathcal{G}$ is an I-map of $e_{f}({\mathbf{x}})$). As the measure induced by $e_{f}({\mathbf{x}})$ is strictly positive, Hammersley-Clifford Theorem clifford1971markov implies that $e_{f}({\mathbf{x}})$ factorizes according to $\mathcal{G}$. That is, for the clique set $\mathcal{C}$ of $\mathcal{G}$, there exist positive functions $\{u_{C}({\mathbf{x}}_{C})\ |\ C\in\mathcal{C}\}$, such that
@@ B.3 Proof of Theorem 2
Following the standard literature of regression agarwal2019reinforcement and using Assumption (3), by leveraging $f_{C}\in\mathcal{F}_{C}$, we can obtain with probability $1-\delta$,
@@
$\geq\sum_{C}\mathbb{E}_{{\mathbf{x}}\sim p}\left[\{\hat{f}_{C}({\mathbf{x}}_{C})-f_{C}({\mathbf{x}}_{C})\}^{2}\right]$
$+\sigma\times\left\{\mathbb{E}_{{\mathbf{x}}\sim p}\left[\{\hat{f}_{C_{1}}({\mathbf{x}}_{C_{1}})-f_{C_{1}}({\mathbf{x}}_{C_{1}})\}^{2}\right]^{1/2}\right\}\times\left\{\mathbb{E}_{{\mathbf{x}}\sim p}\left[\left\{\hat{f}_{C_{2}}({\mathbf{x}}_{C_{2}})-f_{C_{2}}({\mathbf{x}}_{C_{2}})\right\}^{2}\right]^{1/2}\right\}$ (11)